Find the integral surface of the linear PDE
x(y²+z)p - y(x²+z)q = (x²-y²)z , which contains the straight line x+y =0 , z=1 .
1
Expert's answer
2019-08-30T09:57:10-0400

Auxiliary equations are


"{dx\over x(y^2+z)}={dy \over -y(x^2+z)}={dz \over z(x^2-y^2)}"

By Choosing multipliers "x, y, -1," we get


"{xdx+ydy-dz\over x^2y^2+x^2z-x^2y^2-y^2z-x^2z+y^2z}={xdx+ydy-dz\over 0}"

Then


"x^2+y^2-2z=C_1"

By Choosing multipliers "1\/x, 1\/y, 1\/z," we get


"{{dx \over x}+{dx \over x}+{dz \over z}\over y^2+z-x^2-z+x^2-y^2}={{dx \over x}+{dx \over x}+{dz \over z}\over 0}"

Then


"\ln(xyz)=\ln (C_2)"

Or


"xyz=C_2"

Parametric equation of straight line is


"x=t, y=-t, z=1"

Substitute


"t^2+(-t)^2-2(1)=C_1""t(-t)(1)=C_2"

Eliminate "t"


"2t^2-2=C_1""t^2=-C_2"

Then


"-2C_2-2=C_1"

Or


"C_1+2C_2+2=0"

Hence, the  integral surface, which contains the straight line


"x^2+y^2-2z+2xyz+2=0"